# Ultrafast dynamics of the Rydberg states of CO2: Autoionization and Dissociation lifetimes

### v1

@@ III.2 Angular distribution
Figure 5: Angular dependence of differential photoionization and autoionization signal extracted from centered and symmetrized VMI images. (a) Differential photoelectron VMI image for positive time delay showing three red photoelectron rings for $nd$ Henning sharp states and autoionization electrons in blue, (b) Angular distribution of 4d, 5d and 6d photoionization electrons at 250 fs time delay.(b) Angular distribution of autoionization depletion signals integrated from 100-250 fs In each figure, the vertical purple arrow corresponds to the polarization direction for both XUV and NIR pulses.
@@ III.3 Modeling predissociation and autoionization
To understand the role of autoionization and predissociation dynamics we used a model [4] for time-delay dependent photoelectron yield taking both decay channels into account. Our XUV pulse excites neutral ground state population to an excited molecular Rydberg state. This population can decay via autoionization with a rate of $\gamma_{A}$ or via predissociation with a rate of $\gamma_{D}$. If initial excited state population is $P_{0}$, time dependent population in the molecular Rydberg state can be written as
@@
Delayed NIR probe can photoionize the population available in either molecular Rydberg state or in predissociated neutral fragments. Both the channels will produce photoelectrons of same kinetic energies. Thus total population available for NIR photoionization at time t can be written as
@@
Total photoionization rate with a Gaussian NIR pulse at time delay $\tau$ is
@@
where $\sigma$ is the width of the width of the Gaussian pulse. By integrating over time t we get the photoelectron yield as a function of pump-probe time delay.
@@ III.4 Autoionization and predissociation lifetimes
Using the model above we fit experimentally observed photoelectron yields for 4d, 5d and 6d Rydberg states to extract autoionization and predissociation lifetimes. In the discussion of the experimental spectrogram Fig. 3 we mentioned that there is contribution to the signals from the pump-probe cross-correlation around time delay zero. To take that into account, we added another Gaussian contribution to the photoelectron yield given in Eq. 6.
As a first step, we use this Gaussian term to fit the photoelectron yield in the NIR dressing sideband near 1.6 eV (Fig. 6(a)). Fitting the XUV-NIR cross-correlation allows us to extract the width of the probe pulse while also calibrating the zero delay. The standard deviation ($\sigma$) width of the Gaussian cross-correlation is obtained to be $29.77$ fs with an estimated error of $3\%$, while the correction for zero time delay between pump and probe pulse is obtained within $\sim 1$ fs accuracy. Using these parameters of the NIR pulse we fit three Rydberg photoelectron yields as shown in Fig. 6(b)-(d). In each plot, green line represents the XUV-IR cross-correlation signal, blue line corresponds to the Rydberg state dynamics, while the red line is a fit to the total signal. Using this method, we extracted the autoionization and predissociation lifetimes as listed below in Table 1.
Table 1: Autoionization ($\tau_{A}$) and predissociation ($\tau_{D}$) lifetimes of 4d, 5d and 6d Henning sharp series. Numbers in the parenthesis account for the range of fitting parameters for 95% confidence interval.
@@
4d 285.92 (265.04, 310.37) 1782.94 (1536.09, 2124.49)
5d 338.97 (309.88, 374.11) 2574.04 (2002.80, 3652.30)
6d 426.83 (387.89, 474.38) 3054.73 (2079.00, 5757.05)
There is no experimental measurements for the lifetimes of these states available in the literature. Recently, autoionization lifetimes were calculated for these states in Fidler et al [26]. Our experimental lifetime for 5d sharp series matches well within the error limits with the calculated value of 372.6 fs, whereas the lifetime of 6d state does not match with the calculated value of 1457.5 fs. We suspect that the complexity of involving the predissociation process with autoionization causes this discrepancy between the calculation and the experiment. In the experiment by Fidler et al [26], they observe only Henning diffuse state lifetimes due to two-photon Raman transitions with nearby dark resonances. Apart from the experimental measurments of the autoionization lifetimes of Henning sharp states, we have also obtained the experimental predissociation lifetimes of these states. Although there is a significantly large error bar on the predissociation lifetimes, there is no calculation or experimental measurement of these lifetimes available in literature.
@@ IV Conclusion
In this work we used ultrafast time-resolved photoelectron spectroscopy to study the Rydberg state dynamics of $CO_{2}$ molecule. Using delay-dependent photoelectron kinetic energy and angular distribution measurements we identified strong signals for three Henning sharp Rydberg states (4d, 5d and 6d). Using a model that accounts for the competition between autoionization and predissociation, we report the first experimental measurements of both autoionization and predissociation lifetimes for these three Henning sharp states. This XUV-NIR photoelectron spectroscopy method can be generalized to observe the non-adiabatic relaxation dynamics associated with electronic correlations and electron-nuclear couplings in a variety of polyatomic molecules. The knowledge of relaxation rates and dynamics of excited molecules has implications for understanding the charge and energy redistribution in a variety of systems.

### v2

@@ III.2 Angular distribution
Figure 5: Angular dependence of differential photoionization and autoionization signal extracted from centered and symmetrized VMI images. (a) Differential photoelectron VMI image for positive time delay showing three red photoelectron rings for $nd$ Henning sharp states and autoionization electrons in blue, (b) Angular distribution of 4d, 5d, and 6d photoionization electrons at 250 fs time delay.(b) Angular distribution of autoionization depletion signals integrated from 100-250 fs In each figure, the vertical purple arrow corresponds to the polarization direction for both XUV and NIR pulses.
@@ III.3 Modeling predissociation and autoionization
To understand the role of autoionization and predissociation dynamics, we used a model [4] for time-delay dependent photoelectron yield taking both decay channels into account. Our XUV pulse excites neutral ground state population to an excited molecular Rydberg state. This population can decay via autoionization with a rate of $\gamma_{A}$ and via predissociation with a rate of $\gamma_{D}$. If initial excited state population is $P_{0}$, time-dependent population in the molecular Rydberg state can be written as
@@
Delayed NIR probe can photoionize the population available in either molecular Rydberg state or in predissociated neutral fragments. Both channels will produce photoelectrons of same kinetic energies. Thus, the total population available for NIR photoionization at time t can be written as
@@
The total photoionization rate with a Gaussian NIR pulse at time delay $\tau$ is
@@
where $\sigma$ is the width of the Gaussian pulse. By integrating over time t we get the photoelectron yield as a function of pump-probe time delay.
@@ III.4 Autoionization and predissociation lifetimes
Using the model above we fit experimentally-observed photoelectron yields for 4d, 5d, and 6d Rydberg states to extract autoionization and predissociation lifetimes. In the discussion of the experimental spectrogram Fig. 3, we mentioned that there is a contribution to the signal from the pump-probe cross-correlation around time delay zero. To take that into account, we added another Gaussian contribution to the photoelectron yield given in Eq. 6.
As a first step, we use this Gaussian term to fit the photoelectron yield in the NIR dressing sideband near 1.6 eV (Fig. 6(a)). Fitting the XUV-NIR cross-correlation allows us to extract the width of the probe pulse while also calibrating the zero delay. The standard deviation ($\sigma$) width of the Gaussian cross-correlation is obtained to be $29.77$ fs with an estimated error of $3\%$, while the correction for zero time delay between pump and probe pulse is obtained within $\sim 1$ fs accuracy. Using these parameters of the NIR pulse, we fit three Rydberg photoelectron yields as shown in Fig. 6(b)-(d). In each plot, green line represents the XUV-IR cross-correlation signal, blue line corresponds to the Rydberg state dynamics, while the red line is a fit to the total signal. Using this method, we extracted the autoionization and predissociation lifetimes as listed below in Table 1.
Table 1: Autoionization ($\tau_{A}$) and predissociation ($\tau_{D}$) lifetimes of 4d, 5d, and 6d Henning sharp series. Numbers in the parenthesis account for the range of fitting parameters for 95% confidence interval.
@@
4d 286 (265, 310) 1783 (1536, 2124)
5d 339 (310, 374) 2574 (2003, 3652)
6d 427 (388, 474) 3055 (2079, 5757)
There is no experimental measurements for the lifetimes of these states available in the literature. Recently, autoionization lifetimes were calculated for these states in Fidler et al [26]. Our experimental lifetime for 5d sharp series matches well within the error limits with the calculated value of 372.6 fs, whereas the lifetime of 6d state does not match with the calculated value of 1457.5 fs. We suspect that the complexity of involving the predissociation process with autoionization causes this discrepancy between the calculation and the experiment. In the experiment by Fidler et al [26], they observe only Henning diffuse state lifetimes due to two-photon Raman transitions with nearby dark resonances. Apart from the experimental measurments of the autoionization lifetimes of Henning sharp states, we have also obtained the experimental predissociation lifetimes of these states. Although there is a significantly-large error bar on the predissociation lifetimes, there is no calculation or experimental measurement of these lifetimes available in literature.
@@ IV Conclusion
In this work we used ultrafast time-resolved photoelectron spectroscopy to study the Rydberg state dynamics of CO2 molecule. Using delay-dependent photoelectron kinetic energy and angular distribution measurements we identified strong signals for three Henning sharp Rydberg states (4d, 5d, and 6d). Using a model that accounts for the competition between autoionization and predissociation, we report the first experimental measurements of both autoionization and predissociation lifetimes for these three Henning sharp states. This XUV-NIR photoelectron spectroscopy method can be generalized to observe the non-adiabatic relaxation dynamics associated with electronic correlations and electron-nuclear couplings in a variety of polyatomic molecules. The knowledge of relaxation rates and dynamics of excited molecules has implications for understanding the charge and energy redistributions in a variety of systems.